### v1

# The causal structure of galactic astrophysics

Abstract
[…]
keywords:
methods: data analysis – methods: statistical – galaxies: formation – galaxies: fundamental parameters – galaxies: statistics
pubyear: 2025pagerange: The causal structure of galactic astrophysicsLABEL:LastPage
1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone is insufficient for determining causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are. Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences.
This paper seeks to develop causal discovery to the point where it can be applied to the entire low-redshift galaxy population, as a method for adding value to traditional correlation or machine learning analyses. This offers the promise of determining whether the empirical links between physical variables reflect causal pathways (indicating a physical operation of one quantity on another) or merely statistical co-variation (indicating an accidental correlation reflecting a physical law in operation elsewhere). This is precisely the kind of information predicted by physical theories, and hence provides great potential for improving constraints on them.
A possibly complete list of previous applications of causal discovery to astrophysics follows. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024); Jin et al. (2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Finally, Jin et al. (2025a) apply a linear causal discovery model to 100 simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve very small numbers of galaxies far from representative of the population at large, and have been restricted to very specific variable sets and scientific questions.
After describing causal discovery and identifying a method suitable for astrophysical problems, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding latent variables. We show explicitly how this helps to pinpoint the physical mechanisms governing the data, which are crucially directional, in addition simply to inducing correlations.
2 Observational data
We base our analysis on the NSA v1.0.1 (Blanton et al., 2011),111www.sdss4.org/dr17/manga/manga-target-selection/nsa/ a value-added catalogue of nearby galaxies that includes inferred quantities (such as stellar mass and star formation rate) in addition to raw observables. It is based primarily on Sloan Digital Sky Survey (SDSS) imaging but employs reprocessed reductions with improved sky subtraction and photometry tailored for extended sources. The catalogue includes galaxies with redshifts $z\lesssim 0.15$, and provides homogenised multi-band photometry and spectroscopic redshifts, including a cross-match with Galaxy Evolution Explorer (GALEX) data to fill in the ultra-violet part of galaxies’ spectral energy distributions. The NSA is a widely-studied standard for local galaxies (e.g. Reines et al. 2013; Wheeler et al. 2014; Ma et al. 2014; Bundy et al. 2015; Latimer et al. 2021).
We take the following fields:
  • zdist: estimated cosmological redshift correcting the observed redshift with a peculiar velocity estimate. This corresponds approximately to distance in Mpc/$h$.
  • elpetro_absmag: absolute magnitude (luminosity) in the SDSS $r$-band.
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myrs.
  • elpetro_mass: Stellar mass in $M_{\odot}/h^{2}$.
  • sersic_n: Sérsic index $n$ from a two-dimensional, single-component Sérsic fit in the $r$-band. This indicates morphology, with $n=1$ describing an exponential disk (extremely late-type) and $n=4$ a de Vaucouleurs profile (early-type).
  • elpetro_ba: Axis ratio $b/a$ at the isophotal contour enclosing 90 per cent of a galaxy’s light. This also indicates morphology, although affected by projection effects differently to $n$: low $b/a$ indicates a thin, edge-on disk, while high $b/a$ indicates a spheroid or highly inclined galaxy.
  • elpetro_th50_r: Angular radius enclosing 50 per cent of a galaxy’s light in the $r$-band, in arcseconds. (This could be converted to a more physically meaningful absolute size, but we do not do so for this pathfinder study because the causal link that must exist between redshift and apparent size will act as a check on the method.)
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton & Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logs have base 10.
To clean the catalogue we require zdist <0.15<0.15, elpetro_absmag <10<-10, elpetro_b300 $10^{-8}$, elpetro_b300 <10<10, elpetro_mass $10^{6}$, elpetro_mass $10^{12}$, sersic_n <6<6, elpetro_ba >0>0, elpetro_ba <1<1, elpetro_th50_r >0>0 and elpetro_th50_r <25<25. These cuts remove anomalous objects whose properties are likely to have been inaccurately determined. This leaves 587,338 out of an original 641,409 galaxies. A corner plot of the final dataset is shown in Fig. 1.
[…]
3 Methodology
3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. These predictions can be projected onto the space of correlations, but information and hence constraining power is lost in doing so. Causal discovery methods seek retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2014); Eberhardt (2017); Glymour et al. (2019).
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this is possible, suppose we measure three galaxy properties: stellar mass ($M$), gas mass ($G$), and star-formation rate ($S$). Observationally we find that all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously building stellar mass, $S\rightarrow M$) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ remain correlated even after conditioning on $G$, there must exist a direct causal link between them ($M\rightarrow S$); conversely if conditioning on $G$ removes the MSM-S correlation we must instead have $M\rightarrow G\rightarrow S$. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs. Each such class contains the set of causal structures with the same statistical dependencies, and which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: for example $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm, which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible. [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. Alternative score-based methods such Greedy Equivalent Search assign likelihood scores to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
Causal structure is visualised in a causal graph—a directed network of relationships among the variables. The true, generating causal structure is described by a directed acyclic graph (DAG), in which all edges (i.e. causal links between variables, represented by lines) are directed. The corresponding Markov equivalence class is shown by a completed partially directed acyclic graph (CPDAG), which contains edges for which directionality cannot be established based solely on the conditional independencies. A partial ancestral graph (PAG) additionally models the potential effect of latent confounding variables. These can create dependencies between observables that cannot be resolved by any orientation of arrows among the observed variables alone: every candidate orientation produces an inconsistency with other independencies. This replaces the DAG with a maximal ancestral graph (MAG), indicating the possible influence of hidden causes with a circle endpoint. A PAG then represents the Markov equivalence class of the true MAG, and is the goal of algorithms that do not assume causal sufficiency.
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly its own cause). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to weed out insignificant correlations.
3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear correlations (see Fig. 1), and 3) efficient enough to analyse $\mathcal{O}(10^{5})$ objects in reasonable time.
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al 2025, in prep) as implemented in py-tetrad (Ramsey & Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm which starts with a score-based estimate of the causal graph and then identifies a minimal set of conditional independence tests required to remove a causal link between two variables, given the provisional PAG. This reduces the number of tests required and the statistical noise they introduce. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence. The resulting graphs are edge-minimal, correctly oriented and exhibit high accuracy on causal discovery benchmarks, with a runtime of only $\sim$1 minute for $5\times 10^{5}$ datapoints and 7 features.
For independence test we use use_basis_function_lrt, which expands variables into a nonlinear basis set and performs a likelihood ratio test between different conditional independence models. For scoring we adopt use_basis_function_bic, which uses the same nonlinear expansion but employs a Bayesian information criterion (BIC) score with a custom penalty weight:
$\text{BIC}=\mathcal{L}-\texttt{penalty\_discount}\times k\ln(N),$ (1)
where $\mathcal{L}$ is the likelihood of a target variable when predicted according to a hypothetical conditional correlation structure, $N$ is the sample size used for the local regression, and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set.
The two main hyperparameters affecting dataset-specific performance are penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised (Eq. 1). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit. The latter controls the complexity of the local regression model through the number of polynomial basis terms included. Larger values allow more expressive models at the cost of runtime and reduced BIC.
3.3 Mock data generation
To optimise these hyperparameters for our astrophysical application we create mock datasets with similar characteristics to the real data but with known causal structure. This will also enable the reliability of the method to be quantified.
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al 2025, in prep), a code for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions, CPN uses randomly configured multilayer perceptrons. Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. This lets CPN produce highly flexible data approximating a broad class of nonlinear functions.
We generate datasets with the same size as the NSA subset (587,338 points) with 7 nodes and a random number of edges between 12 and 16, roughly matching what will be measured in the real data. This takes around a minute per dataset. We use four hidden layers with 50 neurons each, a ReLU activation function and the default noise distribution $B(2,5)$. This produces mock datasets with correlations visually similar to Fig. 1, but we also check that the results are not sensitive to reasonable variations in them. We then refit each of these datasets with the FCIT algorithm for a range of truncation_limit and penalty_discount values. For each one we compute the precision (fraction of predicted edges that are correct), recall (fraction of true edges that were successfully recovered) and F1 score (harmonic mean of the precision and recall) of the PAG produced.
4 Results
4.1 Mock data tests
We find that truncation_limit =14=14 is ideal for this data: it is considerably larger than the default value of 3, allowing the highly nonlinear relations between variables to be captured, but still larger values tend to decrease the BIC due to the additional model complexity. The results are largely insensitive to this. Fixing this we then scan through penalty_discount, calculating in each case the average precision, recall and F1 score across 200 mock datasets differing only in their number of edges and the random number generation.
The result is shown in Fig. 2. As penalty_discount increases, the scoring function penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but it also causes some true edges to be missed, lowering recall. The F1 score exhibits a peak at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data, but again explore reasonable variations without finding important differences. We also confirm that the other hyperparameters in the FCIT algorithm and its testing and scoring methods have little impact on the results.
[…]
4.2 Real data
We now apply the FCIT algorithm to the NSA data. The PAG produced is shown in Fig. 3.
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift influences the apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance. Mass is seen to causally determine size (rather than vice versa, as would be possible given simply the mass–size relation), suggesting inside-out growth of discs and size expansion via mergers. It also determines Sérsic index, agreeing with the idea that bulge growth and morphological transformation are primarily consequences of hierarchical mass assembly. The absence of link from morphology to mass disfavours simplistic models where concentration alone sets stellar mass. The link from SFR to absolute magnitude reflects the brightening of galaxies in optical bands due to recent star formation.
The uncertain edges between star formation, stellar mass, and morphology highlight the complexity of baryonic processes. The ambiguous edge between stellar mass and luminosity is unsurprising, since mass estimates are derived from photometry and strongly depend on mass-to-light ratios. The circle endpoints highlight the possible role of latent factors not included in the analysis—such as dust attenuation, gas content, and halo environment—which can drive correlations and obscure true directions. The graph does not unambiguously support a picture in which star formation determines morphology on short timescales, or that mass quenching is the sole pathway. It does however imply that the backbone of galaxy evolution—mass driving size and morphology, and star formation driving luminosity—is recoverable directly from survey data. This is highly promising for future, more sophisticated applications of the methodology.
Figure 3: The PAG of the NSA data. Confident causal structures are indicated by directed edges, while less confident associations (circle endpoints) may be impacted by latent confounders.
5 Discussion and Conclusion
The application of causal discovery to astrophysics is largely virgin territory. By enabling the directions of physical links to be established, it provides a significant information overlay on (even machine learning-based) correlation analyses, helping to constrain theories which postulate the physical mechanisms governing the data. Such theories essentially correspond to DAGs, so causal discovery can be seen as a method for inferring theories (as far as is possible) directly from data.
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 3. This supports a hierarchical and mass-driven framework of galaxy evolution while indicating the complexities in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics.
In the near term there are several ways in which this analysis could be extended. First, many of the causal links in Fig. 3 reflect observational or selection effects rather than physical mechanisms. The data could be finessed to minimise these, for example by conditioning properties on redshift or constructing combinations of variables less prone to selection biases or trivial correlations. Second, the several ambiguous (circle) endpoints indicate the potential impact of latent variables not included in the dataset. By folding in such properties as gas mass, metallicity, dust attenuation and environmental density these ambiguities could be resolved, providing a clearer picture of the overall flow of causality. There is of course a huge range of further data across astrophysics that could profitably be interpreted through a causal discovery lens.
There is room for improvement on the theoretical side too. While we showed good performance, our method still relies on choices of conditional independence tests and scoring which have not been explored exhaustively. This could be investigated by applying causal discovery to cosmological simulations, which have known physical mechanisms but more accurately capture likely astrophysical correlations. This would also provide a platform for investigating selection effects in detail, as well as revealing more clearly the causal graphs associated with candidate physical models. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity or non-Gaussianity), as has already been done when assuming causal sufficiency (e.g. Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2023; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery.
In summary, this study paves the way for causal discovery to become as mainstream in astrophysics as it is in other data-rich fields where causal correlations—and their directions—encode crucial information about the underlying mechanisms.
Data availability
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code and all other data will be made available on reasonable request to the corresponding author.
Acknowledgements
We thank David Bacon, Deaglan Bartlett, Robin Evans, Pedro Ferreira, Sebastian von Hausegger, Matt Jarvis, Richard Stiskalek and Tariq Yasin for useful discussions. HD is supported by a Royal Society University Research Fellowship (grant no. 211046).