### v2

# The causal structure of galactic astrophysics

Abstract
[…]
1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone cannot determine causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are. Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences.
This paper seeks to develop causal discovery to the point where it can be applied to the entire low-redshift galaxy population, as a method for adding value to traditional correlation or machine learning analyses. This offers the promise of determining whether the empirical links between physical variables reflect causal pathways (indicating a physical operation of one quantity on another) or merely statistical co-variation (indicating an accidental correlation reflecting a physical law in operation elsewhere). This is precisely the kind of information predicted by physical theories, and hence provides great potential for improving constraints on them.
A possibly complete list of previous applications of causal discovery to astrophysics follows. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024); Jin et al. (2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Davis et al. (2025) use Fast Causal Inference to help constrain the formation histories of trans-Neptunian objects. Finally, Jin et al. (2025a) apply a linear causal discovery model to 100 simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve samples too small to be representative of the underlying populations, and have been restricted to very specific variable sets and scientific questions.
After describing causal discovery and identifying a method suitable for astrophysical problems involving hundreds of thousands of objects, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding variables not in the data subset. We show explicitly how this can be used to refine our understanding of galaxy evolution and test proposed physical mechanisms, which crucially (as is typical in physics) come with a direction of operation in addition to simply inducing correlation among galaxy properties.
2 Observational data
We base our analysis on the NSA v1.0.1 (Blanton et al., 2011),111www.sdss4.org/dr17/manga/manga-target-selection/nsa/ a value-added catalogue of nearby galaxies that includes inferred quantities (such as stellar mass and star formation rate) in addition to raw observables. It is based primarily on Sloan Digital Sky Survey (SDSS) imaging but employs reprocessed reductions with improved sky subtraction and photometry tailored for extended low-redshift sources. The catalogue includes galaxies with redshifts $z\lesssim 0.15$, and provides homogenised multi-band photometry and spectroscopic redshifts, including a cross-match with Galaxy Evolution Explorer (GALEX) data to fill in the ultra-violet part of galaxies’ spectral energy distributions. The NSA is a widely-studied standard for the local galaxy population (e.g. Reines et al. 2013; Ma et al. 2014; Wheeler et al. 2014; Bundy et al. 2015; Latimer et al. 2021).
We take the following fields:
  • zdist: estimated cosmological redshift correcting the observed redshift with a peculiar velocity estimate from Willick et al. (1997). This corresponds approximately to distance in Mpc/$h$.
  • elpetro_absmag: absolute magnitude (luminosity) in the SDSS $r$-band.
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myr.
  • elpetro_mass: Stellar mass in $M_{\odot}/h^{2}$.
  • sersic_n: Sérsic index $n$ from a two-dimensional, single-component Sérsic fit in the $r$-band. This indicates morphology, with $n=1$ describing an exponential disk (extremely late-type) and $n=4$ a de Vaucouleurs profile (early-type).
  • elpetro_ba: Axis ratio $b/a$ at the isophotal contour enclosing 90 per cent of a galaxy’s light. This also indicates morphology, although affected by projection effects differently to $n$: low $b/a$ indicates a thin, edge-on disk, while high $b/a$ indicates a spheroid or highly inclined galaxy.
  • elpetro_th50_r: Angular radius enclosing 50 per cent of a galaxy’s light in the $r$-band, in arcseconds. (This could be converted to a more physically meaningful absolute size, but we do not do so for this pathfinder study because the causal link that must exist between redshift and apparent size will act as a check on the method.)
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. Note that they are by no means unique: different choices would reflect different user priorities and beliefs about the most causally relevant variables. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton & Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logarithms are base 10.
To clean the catalogue we require zdist <0.15<0.15, elpetro_absmag <10<-10, elpetro_b300 $10^{-8}$, elpetro_b300 <10<10, elpetro_mass $10^{6}$, elpetro_mass $10^{12}$, sersic_n <6<6, elpetro_ba >0>0, elpetro_ba <1<1, elpetro_th50_r >0>0 and elpetro_th50_r <25<25. These cuts remove anomalous objects whose properties are likely to have been inaccurately determined. This leaves 587,338 out of an original 641,409 galaxies. A corner plot of the final dataset is shown in Fig. 1.
[…]
3 Methodology
3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. Projecting these predictions onto the space of correlations loses information and hence constraining power. Causal discovery methods seek to retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. It is this information that connects directly to theories and models of galaxy formation and evolution. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2014); Eberhardt (2017); Glymour et al. (2019).
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this works, suppose we measure three galaxy properties: stellar mass ($M$), gas mass ($G$), and star-formation rate ($S$). We find all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously building stellar mass, $S\rightarrow M$) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ remain correlated even after conditioning on $G$, there must exist a direct causal link between them ($M\rightarrow S$), while if conditioning on $G$ removes the MSM-S correlation we must instead have $M\rightarrow G\rightarrow S$. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs. Each class comprises causal structures sharing the same statistical dependencies which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: e.g. $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm, which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible.  [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. There are two more rules involving collisions that can be used to determine directions. Alternative score-based methods such Greedy Equivalent Search assign likelihoods to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
Causal structure is visualised in a causal graph—a directed network of relationships among the variables. The true, generating causal structure is described by a directed acyclic graph (DAG), in which all edges (i.e. causal links between variables, represented by lines) are directed. The corresponding Markov equivalence class is shown by a completed partially directed acyclic graph (CPDAG), which contains edges for which directionality cannot be established based solely on the conditional independencies. A partial ancestral graph (PAG) additionally models the potential effect of latent confounding variables, overcoming the assumption that all causes are contained within the dataset. These can create dependencies between observables that cannot be resolved by any orientation of arrows among the observed variables alone: every candidate orientation produces an inconsistency with other independencies. This replaces the DAG with a maximal ancestral graph (MAG). A PAG then represents the Markov equivalence class of the true MAG, indicating the possible influence of hidden causes with a circle endpoint. This is the goal of causal discovery.
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly its own cause). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to identify insignificant correlations. Some methods further assume causal sufficiency, namely that all causally relevant variables are included in the dataset. Since causal discovery cannot guarantee the “true” causal graph, but rather identifies the set of structures compatible with the observed data and assumptions, it is best used to weed out statistically implausible causal relationships and hence generate testable causal hypotheses to be compared with theoretical expectations or followed up with targeted observations.
3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear functional relationships (see Fig. 1), and 3) efficient enough to analyse $N\approx\mathcal{O}(10^{5})$ objects in reasonable time. This is impossible for traditional algorithms, which scale as $N^{3}$ or worse when accounting for confounders and nonlinearity and become prohibitively memory intensive with $\gtrsim 500$ samples (Strobl et al., 2017; Raghu et al., 2018).
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al. 2025a) as implemented in py-tetrad (Ramsey & Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm which starts with a score-based estimate of the causal graph, possibly with additional adjacencies that the score-based method cannot remove directly, and then applies a recursive procedure to identify minimal sets of conditional independence tests for removing these extra causal links between variables that should be non-adjacent in the correct PAG. This reduces the number of tests required and the statistical noise they introduce. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence. Additional bookkeeping in the FCIT procedure guarantees that interim and final graphs are guaranteed to be well-formed PAGs when applied to data samples, a departure from existing latent variable procedures that yield PAGs theoretically. The resulting graphs are edge-minimal, correctly oriented and exhibit high accuracy on causal discovery benchmarks, with an unprecedented runtime of only $\sim$1 minute for $5\times 10^{5}$ datapoints and 7 features.
For independence test we use use_basis_function_lrt, which expands variables into a nonlinear basis set and performs a likelihood ratio test between different conditional independence models. For scoring we adopt use_basis_function_bic, which uses the same nonlinear expansion but employs a Bayesian information criterion (BIC) score with a custom penalty weight:
$\text{BIC}=\mathcal{L}-\texttt{penalty\_discount}\times k\ln(N),$ (1)
where $\mathcal{L}$ is the likelihood of a target variable according to a hypothetical conditional correlation structure, $N$ is the sample size used for the local regression, and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set (Ramsey et al., 2025b). This makes FCIT a hybrid algorithm: the constraint phase uses the basis‑function likelihood ratio test to prune edges, while the score phase (and the refinement and orientation phases) uses the basis‑function BIC. For the $p$-value threshold we use 0.01 in all cases, which we find not to affect the results appreciably.
Two main hyperparameters affect performance in a dataset-specific way: penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised (Eq. 1). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit as described by $\mathcal{L}$. The latter controls the complexity of the local regression model through the number of polynomial basis terms included. Larger values allow more expressive models at the cost of runtime and enhanced BIC complexity penalty.
3.3 Mock data generation
To optimise these hyperparameters for our astrophysical application we create mock datasets with similar characteristics to the real data but with known causal structure. This will also enable the reliability of the method to be quantified, giving an indication of the systematic uncertainty possible in the results on the real data.
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al. 2025b), a simulation framework for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions CPN uses randomly configured multilayer perceptrons, ensuring that the resulting data exhibits realistic nonlinear dependencies while remaining stable even with a large number of variables. Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. This lets CPN produce highly flexible data approximating a broad class of nonlinear functions.
We generate datasets with the same size as the NSA subset (587,338 points) with 7 nodes and a random number of edges between 12 and 16, roughly matching what will be measured in the real data. This takes around a minute per dataset. We use four hidden layers with 50 neurons each, a ReLU activation function and the default noise distribution $B(2,5)$. This produces mock datasets with correlations visually similar to Fig. 1, but we also check that the results are not sensitive to reasonable variations in them. We then refit each of these datasets with the FCIT algorithm for a range of truncation_limit and penalty_discount values. For each one we compute the precision (fraction of predicted edges that are correct), recall (fraction of true edges that were successfully recovered) and F1 score (harmonic mean of the precision and recall) of the PAG produced.
4 Results
4.1 Mock data tests
We find that truncation_limit =14=14 is ideal for this data: it is considerably larger than the default value of 3, allowing the highly nonlinear relations between variables to be captured in the FCIT scoring and conditional independence tests, but still larger values tend to decrease the BIC due to the additional model complexity. The results are in any case largely insensitive to this. Fixing this we then scan through penalty_discount, calculating in each case the average precision, recall and F1 score across 200 mock datasets differing only in their number of edges and the random number generation.
The result is shown in Fig. 2. As penalty_discount increases the score penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but also causes some true edges to be missed, lowering recall. The F1 score peaks at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data, but again explore reasonable variations without finding important differences. We also confirm that the other hyperparameters in the FCIT algorithm and its testing and scoring methods impact the results little.
[…]
4.2 Real data
We now apply the FCIT algorithm to the NSA data. The PAG produced is shown in Fig. 3.
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift influences the apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance. Mass is seen to causally determine size (rather than vice versa, as would be possible given simply the mass–size relation), suggesting inside-out growth of discs and size expansion via mergers. It also determines Sérsic index, agreeing with the idea that bulge growth and morphological transformation are primarily consequences of hierarchical mass assembly. The absence of a link from morphology to mass disfavours simplistic models where concentration alone sets stellar mass. The link from SFR to absolute magnitude reflects the brightening of galaxies in optical bands due to recent star formation.
The uncertain edges between star formation, stellar mass and morphology highlight the complexity of baryonic processes. Several appear with circle endpoints, indicating algorithmic uncertainty about direction or the influence of confounding variables. The ambiguous link between stellar mass and luminosity is unsurprising, since mass estimates are derived from photometry and strongly depend on mass-to-light ratios. The circle endpoints highlight the possible role of latent factors not included in the analysis—such as dust attenuation, gas content, and halo environment—which can drive correlations and obscure true directions. The graph does not unambiguously support a picture in which star formation determines morphology on short timescales, or that mass quenching is the sole pathway. Instead, it suggests an intertwined system in which stellar mass remains the fundamental regulator, while star formation and morphology are mutually influenced by common drivers such as gas inflow and environment. Disentangling these physical drivers from observational and selection-induced structure (and external latent variables) requires further work, as indicated directly by the circle edges. The result does however imply that the backbone of galaxy evolution—mass driving size and morphology, and star formation driving luminosity—is recoverable directly from survey data. This is highly promising for future, more sophisticated applications of the methodology.
Figure 3: The PAG of the NSA data. Each node contains a colloquial parameter name in bold as well as the technical variable name in the NSA. Confident causal structures are indicated by directed edges, while less confident associations (circle endpoints) may be impacted by latent confounders.
5 Discussion and Conclusion
The application of causal discovery to astrophysics is largely virgin territory. By enabling the directions of physical links to be established, it provides a significant information overlay on (even machine learning-based) correlation analyses, helping to constrain theories that postulate the physical mechanisms governing the data. Such theories essentially correspond to DAGs, so causal discovery can be seen as a method for inferring theories (as far as is possible) directly from data.
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 3. This supports a hierarchical and mass-driven framework of galaxy evolution while indicating the complexities involved in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics. Such methods are sometimes undervalued in favour of intervention-based methods involving randomised control trials.
In the near term there are several ways in which this analysis could be extended. First, many of the causal links in Fig. 3 reflect observational or selection effects rather than physical mechanisms. The data could be refined to minimise these, for example by conditioning properties on redshift or constructing combinations of variables less prone to selection biases or trivial correlations. Second, the several ambiguous (circle) endpoints indicate the potential impact of latent variables not included in the dataset. By folding in such properties as gas mass, metallicity, dust attenuation and environmental density these ambiguities could be resolved, providing a clearer picture of the overall flow of causality. There is of course a huge range of further data across astrophysics that could profitably be interpreted through a causal discovery lens.
There is room for improvement on the theoretical side too. While we showed good performance, our method still relies on choices of conditional independence tests and scoring which have not been explored exhaustively. Mismatches between the mock and real data could lead to errors being larger in the latter. Besides directly investigating other hyperparameter choices, mock datasets and algorithms, this could be investigated by applying causal discovery to cosmological simulations, which have known physical mechanisms but more accurately capture correlations likely to be present in astrophysical data. This would also provide a platform for investigating the loss of information due to selection effects, reveal more clearly the causal graphs associated with candidate physical models and calibrate how reliably different orientations can be distinguished. In parallel, bootstrap and stability analyses on the observational data could determine which edges are most robust.
General and efficient causal discovery methods remain under active development. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity, non-Gaussianity or heteroskedasticity), as has already been done when assuming causal sufficiency in the linear, non-Gaussian regime (e.g. Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2023; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery. Taken together, these steps would provide a more complete and reliable causal map, ultimately allowing causal discovery to move beyond recovering broad backbones to testing detailed models of how galaxies grow and evolve.
In summary, this study paves the way for causal discovery to become as mainstream in astrophysics as it is in other data-rich fields where causal correlations—and their directions—encode crucial information about the underlying mechanisms.
Data availability
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code used in this paper will be made publicly available on GitHub upon publication.
Acknowledgements
We thank David Bacon, Deaglan Bartlett, Robin Evans, Pedro Ferreira, Sebastian von Hausegger, Matt Jarvis, Richard Stiskalek and Tariq Yasin for useful discussions. HD is supported by a Royal Society University Research Fellowship (grant no. 211046).